# T-Explainer: A Model-Agnostic Explainability Framework Based on Gradients

### v2

@@ Abstract
The development of machine learning applications has increased significantly in recent years, motivated by the remarkable ability of learning-powered systems to discover and generalize intricate patterns hidden in massive datasets. Modern learning models, while powerful, often have a level of complexity that renders them opaque black boxes, resulting in a notable lack of transparency that hinders our ability to decipher their reasoning. Opacity challenges the interpretability and practical application of machine learning, especially in critical domains where understanding the underlying reasons is essential for informed decision-making. Explainable Artificial Intelligence (XAI) rises to address that challenge, unraveling the complexity of black boxes by providing elucidating explanations. Among the various XAI approaches, feature attribution/importance stands out for its capacity to delineate the significance of input features in the prediction process. However, most existing attribution methods have limitations, such as instability, when divergent explanations may result from similar or even the same instance. This work introduces T-Explainer, a novel local additive attribution explainer based on Taylor expansion. It has desirable properties, such as local accuracy and consistency, making T-Explainer stable over multiple runs. We demonstrate T-Explainer’s effectiveness in quantitative benchmark experiments against well-known attribution methods. Additionally, we provide several tools to evaluate and visualize explanations, turning T-Explainer into a comprehensive XAI framework.
@@ 1 Introduction
Artificial Intelligence is not a vision for the future. It is our present reality. Terms such as Neural Networks, Machine Learning, Deep Learning, and other facets of the Artificial Intelligence universe have seamlessly shifted from futuristic concepts to near-ubiquitous elements in our daily discourse. The confluence of recent strides in hardware processing capabilities, abundant data accessibility, and refined optimization algorithms has facilitated the creation of intricate and non-linear machine learning models [16]. In the contemporary landscape, those models have achieved unprecedented performance levels, surpassing what was deemed inconceivable just a few years ago, outperforming human abilities and previously known methods in central research areas [63, 56].
@@
The complex non-linear structures and vast number of parameters inherent in such models pose a challenge to transparent interpretation and comprehension of the rationale behind their decisions. Such characteristic transforms those models into black boxes, wherein one can discern only what inputs are provided and what outputs are produced without a clear understanding of the internal decision-making processes [17]. The absence of transparency gives rise to trust-related apprehensions, hindering the effective deployment of potent machine-learning models in critical applications [2, 8, 43]. Additionally, it poses challenges in adhering to emerging regulatory norms observed in numerous countries [24, 58], thereby complicating compliance efforts.
@@
Widely used machine learning models are impenetrable as far as simple interpretations of their mechanisms go [17]. In this context, Explainable Artificial Intelligence (XAI) has emerged to address the challenges outlined above, seeking to provide human-understandable insights into the complexities of models that are inherently challenging to interpret. Significantly, the field of explainability has progressed with the introduction of innovative methodologies, and research focused on discerning the strengths and limitations of XAI models [2]. Feature attribution/importance methods are particularly relevant, as they are currently the most common explanation type [61, 42]. Such methods aim to quantify the contribution of individual input variables to predictions made by black-box models, providing local insights about the model reasoning [43].
@@
While feature attribution methods prove valuable, they come with drawbacks that diminish the trust and confidence placed in their application. For instance, different methods can result in distinct explanations for the same data instances, making it difficult to decide which outcome to believe [61]. Moreover, many well-known methods, including SHAP [39] and LIME [47], suffer from instability, producing different explanations from different runs on a fixed machine learning model and dataset [22].
@@
To be considered stable, an explanation method must produce consistent explanations in multiple runs on the same and similar instances. Stability is a fundamental objective in XAI, as explainability methods that generate inconsistent explanations for similar instances (or even for the same instance) are challenging to trust. If explanations lack reliability, they become essentially useless. Therefore, to be considered reliable, an explanation model must, at a minimum, exhibit stability [7].
@@
In this work, we introduce T-Explainer, a novel post-hoc explanation method that relies on the solid mathematical foundations of Taylor expansion to perform local and model-agnostic feature importance attributions. T-Explainer is a local additive explanation technique with desirable properties such as local accuracy, missingness, and consistency [39]. Furthermore, T-Explainer is deterministic, guaranteeing stable results across multiple runs and delivering consistent explanations for similar instances.
@@
We conducted several quantitative comparisons against well-known feature attribution methods using controlled synthetic and real-world datasets to evaluate the quality and usefulness of T-Explainer explanations. Our evaluations focused on the consistency, continuity, and correctness properties of the explanations’ content [42]. Since T-Explainer is an additive method [39] such as SHAP, we also performed an additivity preservation comparison. As a result of such an extensive evaluation process, we implemented a suite of quantitative metrics to assess different dimensions of feature importance explainers. These metrics were incorporated into the T-Explainer package, rendering it not just another XAI method but a comprehensive XAI framework.
In summary, the main contributions of this work are:
@@
  • T-Explainer, a stable model-agnostic local additive attribution method derived from the solid mathematical foundation of Taylor expansions. It faithfully approximates the local behavior of black-box models using a deterministic optimization procedure, enabling reliable and trustworthy interpretations.
@@
  • A comprehensive set of comparisons against well-known local feature attribution methods using multiple quantitative metrics, aiming to assess stability, faithfulness, and additivity preservation.
@@
  • A Python framework that integrates T-Explainer with other explainability tools, making the proposed method easy to use and assess in different applications. The T-Explainer, evaluation metrics, datasets, and all related materials are openly available online111 A GitHub link will be provided upon acceptance. .
@@ 2 Related work
Several XAI techniques have been proposed to deal with black-box models, either locally or globally [61]. To contextualize our contributions, we focus the following discussion on post-hoc feature importance/attribution methods [66, 32]. A more comprehensive discussion about XAI can be found in several surveys summarizing existing approaches and their properties [2, 8, 37, 43] and metrics for evaluating explanation methods [67, 4, 42].
@@
Breiman [17] proposed one of the first approaches to identify the features most impacting a model prediction. Breiman’s solution is model-specific (Random Forests) and involves permuting the values of each feature and computing the model loss. Given the feature independence assumption, the method identifies the most important features by prioritizing those that contribute the most to the overall loss.
@@
More general approaches, the so-called model-agnostic techniques, can (theoretically) operate with any machine learning model, regardless of the underlying algorithm or architecture. In this context, LIME (Local Interpretable Model-agnostic Explanations) [47], SHAP (SHapley Additive exPlanations) [39], and their variants [46, 48, 21] are model-agnostic methods widely employed to explain machine learning models’ behavior in healthcare [23], financial market [26], and engineering [33] applications.
@@
Although broadly used, LIME and SHAP have significant drawbacks that demand care when using such techniques. For instance, LIME lacks theoretical guarantees about generating accurate simplified approximations for complex models [2]. Moreover, different simplified models can be fit depending on the random sampling mechanism used by LIME, which can lead to instability to small data perturbation and, sometimes, entirely different explanations by just running the code multiple times [14, 7].
@@
SHAP has a solid theoretical foundation derived from game theory [51] that grants SHAP desirable properties such as local accuracy, missingness, and consistency [39]. However, the exact computation of SHAP values is NP-hard, demanding Monte Carlo sampling-based approximations [61, 43], which introduces instability similar to the LIME method, even in its model-specific variants [40, 18, 19]. To avoid instability, deterministic versions of LIME [68], optimization [34], and learning-based [55] approaches have been proposed, but with the price of increasing the number of parameters to be tuned.
@@
Gradient-based methods are another important family of explanation approaches. Those methods attribute importance to each input feature by analyzing how their changes affect the model’s output, relying on gradient decomposition to quantify those effects. Vanilla Gradient [54] introduced the use of gradients to feature attribution tasks. The method computes partial derivatives at the input point $\mathbf{x}$ with a Taylor expansion around a different point $\mathbf{x}^{\prime}$ and a remainder bias term, for which neither is defined [9]. Monte Carlo sampling can be used to estimate derivatives, but it makes Vanilla Gradient suffer from instability and noise within the gradients.
@@
LRP (Layer-wise Relevance Propagation) [9] identifies properties related to the maximum uncertainty state of predictions by redistributing an importance score back to the model’s input layer [35, 27]. The Integrated Gradients method [60] quantifies the feature importance by integrating gradients from the target input to a baseline instance. Input $\times$ Gradient [52] highlights influential regions in the input space by computing the features’ element-wise product and corresponding gradients from the model’s output. DeepLIFT (Deep Learning Important FeaTures) [53] is based on LRP’s importance scores to measure the difference between the model’s prediction of a target input and baseline instance.
@@
The stability of gradient-based methods depends on factors such as the gradient propagation, the model’s complexity, and the baseline choice [53]. Additionally, most of those methods are designed specifically for explanation tasks in Neural Networks and other models with differentiable parameters, which impairs their application to classifiers such as Random Forests and SVMs.
@@
T-Explainer differs from the methods described above in three main aspects: (i) it is deterministic, T-Explainer defines an optimization procedure based on finite differences to estimate partial derivatives, thus being stable by definition; (ii) T-Explainer relies on just a few hyperparameters, rendering it easy to use; and (iii) T-Explainer is not dependent on baselines. Moreover, the T-Explainer is built upon the solid mathematical foundations of Taylor expansion, which naturally endows it with desirable properties similar to those in SHAP. Although T-Explainer, in theory, is designed to be applied to differentiable models, we show experimentally that it also produces interesting results operating with non-differentiable models, which makes T-Explainer more flexible than previous gradient-based methods.
@@ 3 The T-Explainer
As a real function, $f$ can be linearly approximated through first-order Taylor’s expansion:
@@
The $i$-th gradient element corresponds to the partial derivative of $f$ concerning the $i$-th attribute of $\mathbf{x}$. Note that the gradient of $f$ at $\mathbf{x}$ corresponds to the Jacobian matrix when $f$ is a real-valued function. The dot product between the gradient and the displacement vector $\mathbf{h}$ is a linear map from $\mathbb{R}^{n}$ to $\mathbb{R}$, being well-known as the best linear approximation of $f$ in a small neighborhood of $\mathbf{x}$ [45]. Therefore, the gradient can be used to analyze how small perturbations in the input data affect the model output. The right side of Equation 2 is a linear equation that approximates the behavior of $f$ nearby $\mathbf{x}$, and by being a linear mapping, it is naturally interpretable. In other words, the gradient of a model $f$ can be used to generate explanations.
@@
The explanation model $g_{\mathbf{x}}$ is a local attribution method, i.e., there is a $g_{\mathbf{x}}$ for each $\mathbf{x}$. By definition, the T-Explainer is an additive feature attribution method, as defined by Lundberg and Lee [39], meaning that the importance value attributed to each feature can reconstruct the model prediction by summating those importance values. The coefficient $\phi_{i}$ indicates the feature attribution/importance of the $i$-th attribute to the prediction made by $f$ in $\mathbf{x}$. In T-Explainer, the feature attribution $\phi_{i}$ has a simple and intuitive geometric interpretation, corresponding to the projection of $\nabla f(\mathbf{x})$ on the $i$-th feature axis. Therefore, the more aligned the gradient $\nabla f(\mathbf{x})$ and the $i$-th axis of the feature space, the more important the $i$-th feature is to the decision made by $f$.
@@ 3.1 T-Explainer Properties
According to Lundberg and Lee [39], a “good” explanation method must hold three important properties, namely, Local Accuracy, Missingness, and Consistency. In the following, we show that the T-Explainer approximates Local Accuracy while holding Missingness and Consistency.
@@ 3.1.1 Local Accuracy
The local accuracy property, as defined by Lundberg and Lee [39], states that if $f(\mathbf{x}+\mathbf{z}^{\prime})=g_{\mathbf{x}}(\mathbf{z}^{\prime})=\phi_{0}+\sum_{i=1}^{n}{\phi_{i}{z}^{\prime}_{i}}$, then $g_{\mathbf{x}}$ holds the local accuracy property. The T-Explainer does not exactly satisfy this property but rather approximates it. By construction, we have:
@@
and, from the Taylor expansion remainder theorem [41], there is an upper-bound to the approximation error given by:
@@ 3.1.2 Missingness
The missingness property states that if a feature $x_{i}$ has no impact on the model decision, then $\phi_{i}=0$ (see [39, 59]). In our context, a feature $i$ having no impact in $f$ means that $f$ does not vary (increase or decrease) when only such $i$-th feature is changed (otherwise, the feature would impact the model decision). In other words, $f(x_{1},\dots,x_{i}+{z}^{\prime}_{i},\dots,x_{n})-f(x_{1},\dots,x_{i},\dots,x_{n})=0$, thus, there is no variation in the $i$-th direction and the partial derivative $\phi_{i}=\frac{\partial f(\mathbf{x})}{\partial x_{i}}=0$, ensuring that the T-Explainer holds the missingness property.
@@ 3.1.3 Consistency
Let $f$ and $\tilde{f}$ be two binary classification models. Let’s use the notation $\mathbf{x}^{\prime}\backslash i$ to indicate that the $i$-th feature is disregarded in any perturbation of $\mathbf{x}$ (${z}^{\prime}_{i}=0$ in any perturbation, so ${x}^{\prime}_{i}\backslash i=x_{i}$). An explanation method is consistent if (see Lundberg and Lee [39]), fixing $\mathbf{x}$, $\tilde{f}(\mathbf{x}^{\prime})-\tilde{f}(\mathbf{x}^{\prime}\backslash i)>f(\mathbf{x}^{\prime})-f(\mathbf{x}^{\prime}\backslash i)$ implies $\phi_{i}(\tilde{f})>\phi_{i}(f)$.
Suppose that $\tilde{f}(\mathbf{x}^{\prime})-\tilde{f}(\mathbf{x}^{\prime}\backslash i)>f(\mathbf{x}^{\prime})-f(\mathbf{x}^{\prime}\backslash i)$ holds in a small neighborhood of $\mathbf{x}$, in particular,
@@
According to Lundberg and Lee [39], the Shapley-based explanation is the unique possible additive feature attribution model that (theoretically, see Hooker et al. [30]) satisfies Local Accuracy, Missingness, and Consistency properties. As demonstrated above, the T-Explainer approximates Local Accuracy while holding Missingness and Consistency. Therefore, the T-Explainer is one of the few XAI methods that get close to SHAP regarding theoretical guarantees.

### v3

@@ Abstract
The development of machine learning applications has increased significantly in recent years, motivated by the remarkable ability of learning-powered systems to discover and generalize intricate patterns hidden in massive datasets. Modern learning models, while powerful, often exhibit a complexity level that renders them opaque black boxes, lacking transparency and hindering our understanding of their decision-making processes. Opacity challenges the practical application of machine learning, especially in critical domains requiring informed decisions. Explainable Artificial Intelligence (XAI) addresses that challenge, unraveling the complexity of black boxes by providing explanations. Feature attribution/importance XAI stands out for its ability to delineate the significance of input features in predictions. However, most attribution methods have limitations, such as instability, when divergent explanations result from similar or the same instance. This work introduces T-Explainer, a novel additive attribution explainer based on the Taylor expansion that offers desirable properties such as local accuracy and consistency. We demonstrate T-Explainer’s effectiveness and stability over multiple runs in quantitative benchmark experiments against well-known attribution methods. Additionally, we provide several tools to evaluate and visualize explanations, turning T-Explainer into a comprehensive XAI framework.
@@ 1 Introduction
Artificial Intelligence is not a vision for the future. It is our present reality. Terms such as Neural Networks, Machine Learning, Deep Learning, and other facets of the Artificial Intelligence universe have seamlessly shifted from futuristic concepts to near-ubiquitous elements in our daily discourse. The confluence of recent advances in hardware processing capabilities, abundant data accessibility, and refined optimization algorithms has facilitated the creation of intricate and non-linear machine learning models [12]. In the contemporary landscape, these models have achieved unprecedented performance levels, surpassing what was deemed inconceivable just a few years ago, outperforming human abilities and previously known methods in central research areas [48, 42].
@@
The complex non-linear structures and vast number of parameters inherent in such models pose a challenge to transparent interpretation and comprehension of the rationale behind their decisions. This characteristic transforms those models into black boxes in which one can discern only what inputs are provided and what outputs are produced without a clear understanding of the internal decision-making processes [13]. Lack of transparency causes trust-related apprehensions, hindering the effective deployment of more powerful machine learning models in critical applications [2, 6, 34]. Furthermore, it poses challenges in adhering to emerging regulatory norms observed in numerous countries [16, 44], thereby complicating compliance efforts.
@@
Widely used machine learning models are impenetrable as far as simple interpretations of their mechanisms go [13]. In this context, Explainable Artificial Intelligence (XAI) has emerged to address the challenges outlined above, seeking to provide human-understandable insights into the complexities of inherently challenging models. Significantly, the field of explainability has progressed with the introduction of innovative methodologies and research has focused on discerning the strengths and limitations of XAI models [2].
@@
Feature attribution/importance methods are particularly relevant, as they are currently the most common explanation type [46, 33]. Such methods aim to quantify the contribution of individual input variables to predictions made by black-box models, providing local insights about model reasoning [34]. While feature attribution methods prove valuable, they have drawbacks that diminish trust and confidence in their application. For example, different methods can result in distinct explanations for the same data instances, making it difficult to decide which outcome to believe [46]. Moreover, many well-known methods, including SHAP [30] and LIME [36] suffer instability, producing different explanations from different runs on a fixed machine learning model and dataset [14].
@@
An explanation method must produce consistent explanations in multiple runs on the same and similar instances to be considered stable. Stability is a fundamental objective in XAI, as explainability methods that generate inconsistent explanations for similar instances (or even for the same instance) are challenging to trust. If explanations lack reliability, they are essentially useless. Therefore, to be considered reliable, an explanation model must, at a minimum, exhibit stability [5].
@@
In this work, we introduce T-Explainer, a novel post-hoc explanation method that relies on the solid mathematical foundations of Taylor expansion to perform local and model-agnostic feature importance attributions. T-Explainer is a local additive explanation technique with desirable properties such as local accuracy, missingness, and consistency [30]. Furthermore, T-Explainer is deterministic, guaranteeing stable results across multiple runs and providing consistent explanations for similar instances.
@@
We conducted several quantitative comparisons against well-known feature attribution methods using controlled synthetic and real-world datasets to evaluate the quality and usefulness of the T-Explainer explanations. Our evaluations focused on consistency, continuity, and correctness of the content explanation [33]. As a result of such an extensive evaluation process, we implemented a suite of quantitative metrics to evaluate different dimensions of feature importance explainers. These metrics were incorporated into the T-Explainer package, making it not just another XAI method but a comprehensive XAI framework.
In summary, the main contributions of this work are as follows.
@@
  • T-Explainer, a stable model-agnostic local additive attribution method. It faithfully approximates the local behavior of black-box models using a deterministic optimization procedure, enabling reliable and trustworthy interpretations.
@@
  • A comprehensive set of comparisons against well-known local feature attribution methods using multiple quantitative metrics, aiming to evaluate stability and faithfulness.
@@
  • A Python framework that integrates T-Explainer with other explainability tools, making the proposed method easy to use and assess in different applications. The T-Explainer, evaluation metrics, datasets, and all related materials are available online. [^1] Available at https://github.com/evortigosa/EXplainable_AI
@@ 2 Related work
Several XAI techniques have been proposed to deal with black-box models, either locally or globally [46]. To contextualize our contributions, we focus the following discussion on post-hoc feature importance/attribution methods [51, 23]. A more comprehensive discussion about XAI can be found in several surveys summarizing existing approaches and their properties [2, 6, 28, 34] and metrics to evaluate explanation methods [52, 3, 33].
@@
Breiman [13] proposed one of the first approaches to identify the features that most impact the predictions of a model. Breiman’s solution is model-specific (Random Forests) and involves permuting the values of each feature and computing the model loss. Given the feature independence assumption, the method identifies the most important features by prioritizing those that contribute the most to the overall loss.
@@
More general approaches, the so-called model-agnostic techniques, can (theoretically) operate with any machine learning model, regardless of the underlying algorithm or architecture. In this context, LIME (Local Interpretable Model-agnostic Explanations) [36], SHAP (SHapley Additive exPlanations) [30], and their variants [34] are model-agnostic methods widely employed to explain machine learning models’ behavior in healthcare [15], financial market [18], and engineering [24] applications.
@@
Although widely used, LIME and SHAP have significant drawbacks that require care when using such techniques. For example, LIME lacks theoretical guarantees about generating accurate simplified approximations for complex models [2]. Moreover, different simplified models can be fitted depending on the random sampling mechanism used by LIME, which can lead to instability to small data perturbation and, sometimes, entirely different explanations by just running the code multiple times [5].
@@
SHAP has a solid theoretical foundation derived from game theory that grants SHAP desirable properties such as local accuracy, missingness, and consistency [30]. However, the exact computation of SHAP values is NP-hard, demanding Monte Carlo sampling-based approximations [46, 34], which introduces instability similar to the LIME method, even in model-specific variants [31, 34]. To avoid instability, deterministic versions of LIME [53], optimization [25], and learning-based [41] approaches have been proposed, but at the price of increasing the number of parameters to be tuned.
@@
Gradient-based methods are another important family of explanation approaches. Those methods attribute importance to each input feature by analyzing how their changes affect the model’s output, relying on gradient decomposition to quantify those effects. Vanilla Gradient [40] introduced the use of gradients in feature attribution tasks. The method computes partial derivatives at the input point $\mathbf{x}$ with a Taylor expansion around a different point $\mathbf{x}^{\prime}$ and a remainder bias term, for which neither is defined [7]. Monte Carlo sampling can be used to estimate derivatives, but it makes the Vanilla Gradient suffer from instability and noise within the gradients.
@@
LRP (Layer-wise Relevance Propagation) [7] identifies properties related to the maximum uncertainty state of predictions by redistributing an importance score back to the model’s input layer. The Integrated Gradients method [45] quantifies the feature importance by integrating gradients from the target input to a baseline instance. Input $\times$ Gradient highlights influential regions in the input space by computing the features’ element-wise product and corresponding gradients from the model’s output [34]. DeepLIFT (Deep Learning Important FeaTures) [39] is based on LRP’s importance scores to measure the difference between the model’s prediction of a target input and baseline instance.
@@
The stability of gradient-based methods depends on gradient propagation, the model’s complexity, and the baseline choice [39]. Additionally, most of those methods are designed specifically for explanation tasks in Neural Networks and other models with differentiable parameters.
@@
T-Explainer differs from the methods described above in three main aspects: (i) it is deterministic, T-Explainer defines an optimization procedure based on finite differences to estimate partial derivatives, thus being stable by definition; (ii) T-Explainer relies on just a few hyperparameters, rendering it easy to use; and (iii) T-Explainer is not dependent on baselines. Moreover, T-Explainer is built on the solid mathematical foundations of the Taylor expansion, which naturally endows it with desirable properties similar to those of SHAP. Although T-Explainer, in theory, is designed to be applied to differentiable models, we show experimentally that it also produces interesting results operating with non-differentiable models, which makes T-Explainer more flexible than previous gradient-based methods.
@@ 3 The T-Explainer
As a real function, $f$ can be linearly approximated through first-order Taylor expansion:
@@
The $i$-th gradient element corresponds to the partial derivative of $f$ concerning the $i$-th attribute of $\mathbf{x}$. Note that the gradient of $f$ in $\mathbf{x}$ corresponds to the Jacobian matrix when $f$ is a real-valued function. The dot product between the gradient and the displacement vector $\mathbf{h}$ is a linear map from $\mathbb{R}^{n}$ to $\mathbb{R}$, well known as the best linear approximation of $f$ in a small neighborhood of $\mathbf{x}$.Therefore, the gradient can be used to analyze how small perturbations in the input data affect the model output. The right side of Equation 2 is a linear equation that approximates the behavior of $f$ near $\mathbf{x}$, and because it is a linear mapping, it is naturally interpretable. In other words, the gradient of a model $f$ can be used to generate explanations.
@@
The explanation model $g_{\mathbf{x}}$ is a local attribution method, i.e., there is a $g_{\mathbf{x}}$ for each $\mathbf{x}$. By definition, T-Explainer is an additive feature attribution method, as defined by Lundberg and Lee [30], which means that the importance value attributed to each feature can reconstruct the model prediction by adding those importance values. The coefficient $\phi_{i}$ indicates the feature attribution/importance of the $i$-th attribute to the prediction made by $f$ in $\mathbf{x}$. In T-Explainer, the feature attribution $\phi_{i}$ has a simple and intuitive geometric interpretation, corresponding to the projection of $\nabla f(\mathbf{x})$ on the $i$-th feature axis. Therefore, the more aligned the gradient $\nabla f(\mathbf{x})$ and the $i$-th axis of the feature space, the more important the $i$-th feature is to the decision made by $f$.
@@ 3.1 T-Explainer Properties
According to Lundberg and Lee [30], a “good” explanation method must have three important properties, namely, Local Accuracy, Missingness, and Consistency. In the following, we show that T-Explainer approximates Local Accuracy while holding Missingness and Consistency.
@@ 3.1.1 Local Accuracy
The local accuracy property, as defined by Lundberg and Lee [30], states that if $f(\mathbf{x}+\mathbf{z}^{\prime})=g_{\mathbf{x}}(\mathbf{z}^{\prime})=\phi_{0}+\sum_{i=1}^{n}{\phi_{i}{z}^{\prime}_{i}}$, then $g_{\mathbf{x}}$ has the local accuracy property. The T-Explainer does not fully satisfy this property but instead approximates it. By construction, we have the following:
@@
and, from the Taylor expansion remainder theorem [32], there is an upper bound to the approximation error given by:
@@ 3.1.2 Missingness
The missingness property states that if a feature $x_{i}$ has no impact on the model decision, then $\phi_{i}=0$ [30]. In our context, a feature $i$ having no impact on $f$ means that $f$ does not vary (increase or decrease) when only the $i$-th feature is changed (otherwise, the feature would impact the model decision). In other words, $f(x_{1},\dots,x_{i}+{z}^{\prime}_{i},\dots,x_{n})-f(x_{1},\dots,x_{i},\dots,x_{n})=0$, thus, there is no variation in the $i$-th direction and the partial derivative $\phi_{i}=\frac{\partial f(\mathbf{x})}{\partial x_{i}}=0$, ensuring that the T-Explainer holds the missingness property.
@@ 3.1.3 Consistency
Let $f$ and $\tilde{f}$ be two binary classification models. Let us use the notation $\mathbf{x}^{\prime}\backslash i$ to indicate that the $i$-th feature is disregarded in any perturbation of $\mathbf{x}$ (${z}^{\prime}_{i}=0$ in any perturbation, so ${x}^{\prime}_{i}\backslash i=x_{i}$). An explanation method is consistent if (see Lundberg and Lee [30]), fixing $\mathbf{x}$, $\tilde{f}(\mathbf{x}^{\prime})-\tilde{f}(\mathbf{x}^{\prime}\backslash i)>f(\mathbf{x}^{\prime})-f(\mathbf{x}^{\prime}\backslash i)$ implies $\phi_{i}(\tilde{f})>\phi_{i}(f)$. Suppose that $\tilde{f}(\mathbf{x}^{\prime})-\tilde{f}(\mathbf{x}^{\prime}\backslash i)>f(\mathbf{x}^{\prime})-f(\mathbf{x}^{\prime}\backslash i)$ holds in a small neighborhood of $\mathbf{x}$, in particular,
@@
According to Lundberg and Lee [30], the Shapley-based explanation is the unique possible additive feature attribution model that (theoretically, see Hooker et al. [21]) satisfies Local Accuracy, Missingness, and Consistency properties. As demonstrated above, T-Explainer approximates Local Accuracy while holding Missingness and Consistency. Therefore, the T-Explainer is one of the few XAI methods that approximate SHAP theoretical guarantees.